OpenStudy (anonymous):

solve 3(x+4)-4(x+3)=7x+16 1: x=3 2: x=2 3: x=-2

OpenStudy (texaschic101):

distribute through all of the parenthesis

OpenStudy (anonymous):

man ive been on this question for like 15 min i dont it

OpenStudy (anonymous):

dont get it*

OpenStudy (texaschic101):

3(x + 4) - 4(x + 3) = 7x + 16 3x + 12 - 4x - 12 = 7x + 16 are you with me so far ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

now we will combine like terms -x = 7x + 16 still with me ?

OpenStudy (anonymous):

kpajon are you writing a love letter?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

In this situation you have to input the three solutions. 1. 3(3+4)-4(3+3)=7(3)+16 3(7)-4(6)=21+16 21-24=37 -3 does not equal 37, therefore it is not a solution. 2. 3(2+4)-4(2+3)=7(2)+16 3(6)-4(5)=14+16 18-20=30 -2 does not equal 30, therefore it is not a solution. 3. 3(-2+4)-4(-2+3)=7(-2)+16 3(2)-4(1)=-14+16 6-4=2 2 does equal 2, therefore -2 is the solution.

OpenStudy (anonymous):

lol you better love me after I did all that work -.-

OpenStudy (texaschic101):

-x = 7x + 16 subtract 7x from both sides -x - 7x = 16 -8x = 16 x = -16/8 x = -2 check... 3(x + 4) - 4(x + 3) = 7x + 16 3(-2 + 4) - 4(-2 + 3) = 7(-2) + 16 3(2) - 4(1) = -14 + 16 6 - 4 = -14 + 16 2 = 2 (correct) any questions ?

OpenStudy (anonymous):

ty for the ANSWER!

OpenStudy (anonymous):

lol -2 or 2

OpenStudy (texaschic101):

-2

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

The solution is -2

OpenStudy (texaschic101):

you do not need to plug in all solutions to find the answer....just solve for x.